### v2

# Diagonalization Procedure for Reducible Massless ${\cal N}=1$ Supermultiplets

Diagonalization Procedure for Reducible Massless ${\cal N}=1$ Supermultiplets
Keith Glennon***e-mail: keith.glennon@oist.jp and Mirian Tsulaiae-mail: mirian.tsulaia@oist.jp
Okinawa Institute of Science and Technology,
1919-1 Tancha, Onna-son, Okinawa 904-0495, Japan
We consider ${\cal N}=1$ supersymmetric systems in $d=4$, $6$ and $10$ dimensions which consist of reducible bosonic and fermionic massless representations of the Poincaré group. We show in detail how to decompose the corresponding Lagrangians into a sum of Lagrangians for irreducible representations of the Poincaré group. We also outline a modification of this procedure in the case of an anti-de Sitter background.
1 Introduction
There are various approaches to supersymmetric higher spin theories both on flat and on anti-de Sitter ($AdS$) backgrounds (see [1][5] for a pioneering work in this direction, and [6][10] for reviews). In most cases one considers irreducible representations of the Poincaré or $AdS_{d}$ groups111 A possible implication of ${\cal N}=2$ supersymmetry in higher spin theories on $dS_{4}$ background was recently discussed in [11]., which in turn form irreducible, and in some cases infinite-dimensional, representations of the corresponding supersymmetry algebra. On the other hand, consideration of reducible representations might open some new interesting possibilities. Systems with reducible representations of the Poincaré group were studied in detail in [12][15] and called triplets and generalized triplets, in case fields with mixed symmetries are included. The main advantage of considering reducible representations is that the corresponding free Lagrangians and supersymmetry transformations have a relatively simple form when compared to their irreducible counterparts. This simplicity of the quadratic Lagrangians usually proves to be very helpful when constructing cubic, and higher order, interactions, both for massive and massless higher spin fields [16][19], [20].
Following this line of thought, we elaborate on the ${\cal N}=1$ supersymmetric model in $d=4,6,10$, space-time dimensions, constructed in [21][22]. In this superstring-inspired model, the fermionic sector consists of the fields described by totally symmetric spin–tensors, whereas the bosonic sector consists of both totally symmetric fields and fields with mixed symmetry of the simplest type [^2] Supersymmetry in $d$ dimensions in the unfolded formulation was recently considered in [23]. $Y(n-1,1)$. Because of the gauge invariant description, the system contains, along with physical fields, also fields which are gauge degrees of freedom both in bosonic and in fermionic sectors. One can show that, after the complete fixing of the gauge, the spectrum contains only massless reducible representations of the Poincaré group described by the Young diagrams $Y(n,0)$ and $Y(n-1,1)$ in the bosonic sector and Young diagrams $Y(n-1,0)$ in the fermionic sector. Both bosonic and fermionic representations are reducible, since the traces of the fields are non-zero on-shell. Moreover, the corresponding ${\cal N}=1$ supermultiplets are irreducible in $d=10$, but reducible in $d=4$ and $d=6$ also from the point of view of representations of supersymmetry algebra. As mentioned above, although this approach contains more fields compared to an approach based on irreducible representations of the supersymmetry and Poincaré groups, it allows one to construct Lagrangians for both higher and lower spin fields in a relatively simple form. For example, in this way one can obtain a new description of linearized ${\cal N}=1$ supergravity and linear supermultiplet in four dimensions, linearized ${\cal N}=1$ supergravity and $(1,0)$ tensor multiplet in six dimensions, and $N=1$ linearized supergravity in ten dimensions (see also [24] for a formulation of linearized $d=6$ supergravity in terms of reducible fields).
The purpose of the present paper is to develop a procedure for the decomposition of fermionic and bosonic Lagrangians constructed in [21] into a direct sum of Lagrangians for individual irreducible representations of the Poincaré group [25][26]. This procedure will henceforth be referred to as a diagonalization procedure. For totally symmetric representations of the Poincaré group the diagonalization has been done in [27][29]. Here we show how to perform the diagonalization procedure for fermionic totally symmetric fields and for the fields with the mixed symmetry [^3] See [30][36] for different quadratic actions for mixed symmetry fields on Minkowski background. of the type $Y(n,1)$. Let us note, that without requiring of ${\cal N}=1$ supersymmetry, both bosonic and fermionic Lagrangians as well as the corresponding diagonalization procedures are valid for any ($d\geq 4$) number of space-time dimensions.
The motivation behind developing the diagonalization procedure is two-fold. Firstly, although it is more convenient to construct cubic and higher-order interaction vertices for reducible representations, in certain cases it might still be preferable to consider the interactions between reducible modes. The second reason is the possibility of addressing the question of a modification of the system constructed [21] to an $AdS_{d}$ background [^4] The problem of a formulation of higher spin theories on backgrounds which are different from flat and $(A)dS_{d}$ spaces has been recently addressed in [37].. A deformation of bosonic and fermionic triplets that contain only totally symmetric fields to $AdS_{d}$ [13], [38][39], as well as the relevant diagonalization procedure [28] are known and can be applied directly to the present analysis. However, the presence of mixed-symmetry fields in the spectrum causes some difficulties, and we shall comment on them at the end of Section 4.
The paper is organized as follows. In Section 2 we summarize the main results of [21] which are needed for the purpose of this paper.
In Section 3 we demonstrate in detail the procedure of diagonalization for the simplest case of $n=1$, when the physical fields in the bosonic sector are the second-rank symmetric and antisymmetric tensors, whereas in the fermionic sector one has a first-rank spinor-tensor. We also discuss various ways of deformation of this system to an $AdS_{d}$ background.
In Section 4 we repeat this procedure in the more complicated case of $n=2$. We show that after the diagonalization of the bosonic Lagrangian one obtains the Fronsdal Lagrangian for the spin $3$ field, i.e., for the irreducible representation $Y(3,0)$, a Lagrangian for the mixed symmmetry irreducible representation of type $Y(2,1)$, and two Lagrangians for vector fields, which correspond to the traces of above mentioned (initially reducible) representations. In the fermionic sector one obtains the fields with spins $\frac{5}{2}$, $\frac{3}{2}$ and $\frac{1}{2}$. The detailed study of these examples suggests a clear pattern of how to generalize the procedure to the higher-rank tensors, which is discussed in Section 5.
We conclude with a discussion of some possible applications of our results.
Finally, the Lagrangians for massless irreducible bosonic and fermionic higher spin fields on $AdS_{d}$, and the corresponding algebra of the operators are collected in the Appendix.
2 Reducible ${\cal N}=1$ supermultiplets
Let us briefly summarize the main features of the ${\cal N}=1$ supersymmetric system constructed in [21]. To this end, it is convenient to introduce an auxiliary Fock space, spanned by one bosonic and one fermionic auxiliary oscillators
$[\alpha_{\mu},\alpha^{+}_{\nu}]=\{\psi_{\mu},\psi^{+}_{\nu}\}=\eta_{\mu\nu},\quad\alpha_{\mu}|0\rangle=\psi_{\mu}|0\rangle=0$ (2.1)
Let us consider the bosonic fields first, were one has a physical field $|\phi\rangle$ and a set of auxiliary fields. The fields are of type $|X^{(p,q)}\rangle$, where $p$ is a number of the bosonic oscillators and $q$ which can be either $1$ or 00, is the number of fermionic oscillators. Obviously the fields are symmetrical with respect to $p$-indices and no symmetry between $p$-indices and $q$-index is implied.
In the bosonic sector one has the vectors
$|\phi^{(n,1)}\rangle,\quad|D^{(n-2,1)}\rangle,\quad|A^{(n-1,0)}\rangle,\quad|B^{(n-1,0)}\rangle$ (2.2)
$|C^{(n-1,1)}\rangle,\quad|E^{(n,0)}\rangle,\quad|F^{(n-2,0)}\rangle$ (2.3)
The corresponding quadratic Lagrangian has the form
$-{\cal L}_{bos.}$ $=$ $\langle\phi|l_{0}|\phi\rangle-\langle D|l_{0}|D\rangle+i\langle B|l_{0}|A\rangle-i\langle A|l_{0}|B\rangle$
$+$ $\langle C|l^{+}|D\rangle+i\langle E|l^{+}|B\rangle-\langle C|l|\phi\rangle-i\langle F|l|B\rangle$
$+$ $i\langle C|g^{+}|A\rangle+i\langle E|g|\phi\rangle-i\langle F|g|D\rangle$
$+$ $\langle D|l|C\rangle-i\langle B|l|E\rangle-\langle\phi|l^{+}|C\rangle+i\langle B|l^{+}|F\rangle$
-- $i\langle A|g|C\rangle-i\langle\phi|g^{+}|E\rangle+i\langle D|g^{+}|F\rangle$
$+$ $\langle C|C\rangle+\langle E|E\rangle-\langle F|F\rangle$
with d‘Alembertian, divergence and gradient operators being
$l_{0}=p\cdot p,\quad l^{\pm}=\alpha^{\pm}\cdot p,\quad g^{\pm}=\psi^{\pm}\cdot p,\quad p=-i\partial_{\mu}$ (2.5)
The Lagrangian (2) is invariant under the gauge transformations
$\delta|\phi\rangle=l^{+}|\lambda\rangle+ig^{+}|\rho\rangle,$ (2.6)
$\delta|D\rangle=l|\lambda\rangle+ig^{+}|\sigma\rangle,\quad\delta|A\rangle=-l^{+}|\sigma\rangle+l|\rho\rangle-|\tau\rangle,\quad\delta|B\rangle=-g|\lambda\rangle-i|\tau\rangle,\$
$\delta|C\rangle=l_{0}|\lambda\rangle+ig^{+}|\tau\rangle,\quad\delta|E\rangle=l_{0}|\rho\rangle-l^{+}|\tau\rangle,\quad\delta|F\rangle=l_{0}|\sigma\rangle-l|\tau\rangle$
where the oscillator content for the parameters of gauge transformations is
$|\lambda^{(n-1,1)}\rangle,\quad|\rho^{(n,0)}\rangle,\quad|\tau^{(n-1,0)}\rangle,\quad|\sigma^{(n-2,0)}\rangle$ (2.7)
As one can see from the Lagrangian (2), the fields (2.3) have no kinetic terms and can be expressed via their own equations of motion
$|C\rangle$ $=$ $-l^{+}|D\rangle+l|\phi\rangle-ig^{+}|A\rangle$ (2.8)
$|E\rangle$ $=$ $-il^{+}|B\rangle-ig|\phi\rangle$
$|F\rangle$ $=$ $-il|B\rangle-ig|D\rangle$
Thus the system can be formulated only in terms of the fields (2.2) with the corresponding Lagrangian
${\cal L}_{bos.}$ $=$ $-\langle\phi|l_{0}-l^{+}l-g^{+}g|\phi\rangle+\langle D|2l_{0}+l^{+}l-g^{+}g|D\rangle+\langle M|l_{0}|M\rangle$
-- $\langle\phi|l^{+}l^{+}|D\rangle+\langle\phi|l^{+}g^{+}|M\rangle-\langle D|g^{+}\,l|M\rangle$
-- $\langle D|l\,l|\phi\rangle+\langle M|l\,g|\phi\rangle-\langle M|l^{+}g|D\rangle$
where $|M\rangle=|B\rangle-i|A\rangle$. It can be shown (see [21] for the details), that one can use the gauge invariance (2.6) and the equations of motion in order to eliminate on–shell all fields in (2.2)–(2.3), except $|\phi\rangle$. This vector obeys massless Klein-Gordon equation, is transverse with respect to both sets of its indices and contains the totally symmetric representations of the Poincaré group of type $Y(n+1,0)$, fields with mixed symmetry of type $Y(n,1)$, as well as their traces.
The states in the fermionic sector contain only the bosonic oscillators, defined in (2.1). The fields, as well as a Fock vacuum carry the corresponding $d$dimensional spinorial index, which will not be written explicitly. The field content is
$|\Psi^{(n,0)}\rangle,\quad|\chi^{(n-1,0)}\rangle,\quad|\Sigma^{(n-2,0)}\rangle$ (2.10)
with the corresponding Lagrangian
${\cal L}_{f.}$ $=$ $-\langle\Psi|\gamma\cdot p|\Psi\rangle+\langle\Sigma|\gamma\cdot p|\Sigma\rangle-\langle\Psi|l^{+}|\chi\rangle+\langle\Sigma|l|\chi\rangle$
-- $\langle\chi|l|\Psi\rangle+\langle\chi|l^{+}|\Sigma\rangle-\langle\chi|\gamma\cdot p|\chi\rangle,$
where $\gamma^{\mu}$ are $d$-dimensional gamma matrices, with $\{\gamma^{\mu},\gamma^{\nu}\}=2\,\eta^{\mu\nu}$. The Lagrangian (2) is invariant under gauge transformations
$\delta|\Psi\rangle=l^{+}|\xi\rangle,\quad\delta|\chi\rangle=-\gamma\cdot p|\xi\rangle\quad\delta|\Sigma\rangle=l|\xi\rangle$ (2.12)
where $|\xi^{(n-1,0)}\rangle$ is a parameter. Again, one can show, that it is possible to gauge away the non-physical fields $|\chi\rangle$ and $|\Sigma\rangle$, and one is left on-shell with the physical field $|\Psi\rangle$. This field obeys massless Dirac equation, is transverse and describes a massless field with spin $n+\frac{1}{2}$ as well as lower spins $n-\frac{1}{2},...,\frac{1}{2}$, since no zero gamma–trace condition is implied.
Finally, one can establish ${\cal N}=1$ supersymmetry transformations
$\delta|\Psi\rangle$ $=$ $-(\gamma\cdot p)(\gamma\cdot\psi)\varepsilon|\phi\rangle+i\varepsilon|E\rangle$ (2.13)
$\delta|\Sigma\rangle$ $=$ $-(\gamma\cdot p)(\gamma\cdot\psi)\varepsilon|D\rangle+i\varepsilon|F\rangle$
$\delta|\chi\rangle$ $=$ $(\gamma\cdot\psi)\varepsilon|C\rangle$
$\delta\langle\phi|$ $=$ $\langle\Psi|(\gamma\cdot\psi)\varepsilon$ (2.14)
$\delta\langle D|$ $=$ $\langle\Sigma|(\gamma\cdot\psi)\varepsilon,\quad\delta\langle B|=\langle\chi|\varepsilon$
$\delta\langle C|$ $=$ $\langle\chi|(\gamma\cdot p)(\gamma\cdot\psi)\varepsilon$
which leaves the sum of the Lagrangians (2) and (2) invariant. One can check that the supersymmetry algebra close on-shell in the dimensions $d=4,6$ and $10$. Expressing the auxiliary fields via their own equations of motion (2.8), one can can write the supersymmetry transformation rules for fermions as
$\delta|\Psi\rangle$ $=$ $-\left((\gamma\cdot p)(\gamma\cdot\psi)-g\right)\varepsilon|\phi\rangle+l^{+}\varepsilon|B\rangle$ (2.15)
$\delta|\Sigma\rangle$ $=$ $-\left((\gamma\cdot p)(\gamma\cdot\psi)-g\right)\varepsilon|D\rangle+l\varepsilon|B\rangle$
$\delta|\chi\rangle$ $=$ $l(\gamma\cdot\psi)\varepsilon|\phi\rangle-l^{+}(\gamma\cdot\psi)\varepsilon|D\rangle-i(\gamma\cdot p)\varepsilon|A\rangle$
Let us note that, the closure of the supersymmetry algebra, as well as the consistency of the equations (2.8) with the transformation rules (2.13)– requires the use of fermionic field equations, derived from the Lagrangian (2), as it should have been expected since we are considering on-shell supersymmetry.
3 Diagonalization for $n=1$
3.1 Flat space-time
In order to demonstrate the details of the diagonalization procedure and to see the pattern, we start with the simplest case of $n=1$.
After performing the first step of eliminating auxiliary fields in terms of their own equations of motion, the bosonic Lagrangian (2) reads
${\cal L}_{bos.}=\phi^{\mu,\nu}\Box\phi_{\mu,\nu}-M\Box M+2\partial^{\mu}\phi_{\mu,\nu}\partial^{\nu}M-\phi^{\tau,\nu}\partial_{\tau}\partial^{\mu}\phi_{\mu,\nu}-\phi^{\mu,\tau}\partial_{\tau}\partial^{\nu}\phi_{\mu,\nu}$ (3.1)
The second step is to decompose the field $\phi_{\mu,\nu}$ in terms of the fields with definite symmetries
$\phi_{\mu,\nu}=\tilde{\phi}_{\mu\nu}+{\cal B}_{\mu\nu},\quad\tilde{\phi}_{\nu\mu}=\tilde{\phi}_{\mu\nu},\quad{\cal B}_{\mu\nu}=-{\cal B}_{\nu\mu}$ (3.2)
The Lagrangian (3.1) written in terms of the new fields has the form
${\cal L}_{bos.}={\cal L}_{bos.}^{\prime}+{\cal L}_{bos.}^{\prime\prime}$ (3.3)
where
${\cal L}_{bos.}^{\prime}=\tilde{\phi}^{\mu\nu}\,\Box\,\phi_{\mu\nu}-2\tilde{\phi}^{\mu\tau}\partial_{\tau}\partial^{\nu}\tilde{\phi}_{\mu\nu}+4\tilde{D}\,\partial^{\mu}\partial^{\nu}\,\tilde{\phi}_{\mu\nu}-4\tilde{D}\,\Box\,\tilde{D}$ (3.4)
${\cal L}_{bos.}^{\prime\prime}={\cal B}^{\mu\nu}\,\Box\,{\cal B}_{\mu\nu}-2{\cal B}^{\mu\nu}\,\partial_{\nu}\partial^{\tau}\,{\cal B}_{\mu\tau}$ (3.5)
and we defined $\tilde{D}=-\frac{1}{2}M$. The final step consists of splitting the Lagrangian ${\cal L}_{bos.}^{\prime}$ into a sum of Lagrangians describing fields with spin $2$ and spin 00. An obvious candidate for the gauge invariant scalar is
${\Phi}=\tilde{\phi}^{\mu}{}_{\mu}-2\tilde{D}$ (3.6)
For the spin $2$ field $\Phi_{\mu\nu}$ one takes an ansatz
$\tilde{\phi}_{\mu\nu}=\Phi_{\mu\nu}+\frac{1}{d-2}\eta_{\mu\mu}\Phi$ (3.7)
where the coefficient in front of the second term is determined from the requirement that the Lagrangian ${\cal L}_{bos.}^{\prime}$ splits into a sum of independent Lagrangians for the fields ${\Phi}_{\mu\nu}$ and ${\Phi}$. Indeed, one can check that this choice leads to the Lagrangian
${\cal L}_{bos.}^{\prime}$ $=$ $\Phi^{\mu\nu}\,\Box\,\Phi_{\mu\nu}-2\Phi^{\mu\tau}\partial_{\tau}\partial^{\rho}\Phi_{\mu\rho}-\Phi^{\mu}{}_{\mu}\,\Box\,\Phi^{\nu}{}_{\nu}+2\Phi^{\rho}{}_{\rho}\,\partial_{\mu}\partial_{\nu}\,\Phi^{\mu\nu}$ (3.8)
$+$ $\frac{1}{d-2}\Phi\,\Box\,\Phi$
and this completes the treatment of the bosonic part of the system for the case of $n=1$.
In the fermionic sector according to (2) one has the Lagrangian
${\cal L}_{f.}=i\bar{\Psi}^{\mu}\gamma^{\nu}\partial_{\nu}\Psi_{\mu}+i\bar{\Psi}^{\mu}\partial_{\mu}\chi-i\bar{\chi}\partial^{\mu}\Psi_{\mu}-i\bar{\chi}\gamma^{\nu}\partial_{\nu}\chi$ (3.9)
invariant under the gauge transformations
$\delta\Psi_{\mu}=\partial_{\mu}\xi,\quad\delta\chi=-\gamma^{\mu}\partial_{\mu}\xi$ (3.10)
For the fermionic fields, the first two steps of the diagonalization procedure are not needed and one starts from the third step. Similarly to the case of the bosonic fields we introduce a spin $\frac{3}{2}$ field $\tilde{\Psi}_{\mu}$ and a gauge invariant spin $\frac{1}{2}$ field $\tilde{\Psi}$ as
$\tilde{\Psi}_{\mu}=\Psi_{\mu}+\frac{1}{2-d}\gamma_{\mu}\tilde{\Psi},\quad\tilde{\Psi}=\gamma^{\mu}\Psi_{\mu}+\chi,$ (3.11)
where the coefficient in front of the last term in the definition of the field $\tilde{\Psi}_{\mu}$ is chosen in such a way, that the Lagrangian (3.9) written in terms of the new fields is a sum of Fang – Fronsdal Lagrangians for spins $\frac{3}{2}$ and $\frac{1}{2}$
${\cal L}_{f.}$ $=$ $i\bar{\tilde{\Psi}}^{\mu}\gamma^{\nu}\partial_{\nu}\tilde{\Psi}_{\mu}-i\bar{\tilde{\Psi}}^{\mu}\partial_{\mu}\gamma^{\nu}\tilde{\Psi}_{\nu}-i\bar{\tilde{\Psi}}^{\mu}\gamma_{\mu}\partial_{\nu}\tilde{\Psi}^{\nu}+i(\bar{\tilde{\Psi}}^{\mu}\gamma_{\mu})\gamma^{\nu}\partial_{\nu}(\gamma^{\rho}\tilde{\Psi}_{\rho})$ (3.12)
$+$ $\frac{i}{d-2}\bar{\tilde{\Psi}}\gamma^{\mu}\partial_{\mu}\tilde{\Psi}$
Finally, using the equations (3.6), (3.7), (3.11), as well as (2.14) and (2.15), it is straightforward to rewrite the supersymmetry transformations in terms of irreducible modes. From the field content given above, one can see that this system describes a linearized ${\cal N}=1$ supergravity and the linear multiplet in case of $d=4$, a linearized ${\cal N}=1$ supergravity and $(1,0)$ tensor multiplet in case of $d=6$ and a linearized ${\cal N}=1$ supegravity in case of $d=10$ dimensions [21].
3.2 Deformation to $AdS$ space
In order to perform a deformation of the bosonic (2) and fermionic (2) Lagrangians to an $AdS_{d}$ background, one can proceed in two different ways. The first option is to start with the already decomposed system described by a sum of the Lagrangians (3.5), (3.8) and (3.12). Then we perform its deformation to $AdS_{d}$ space by replacing the usual partial derivatives with the covariant ones and adding the corresponding mass-like terms in order to keep all gauge symmetries intact. The Lagrangians for totally symmetric fermionic and bosonic fields can be directly read off from the equations (A) and (A) which we give here for convenience
${\cal L}_{bos.}^{\prime}$ $=$ $\Phi^{\mu\nu}\,\Box\,\Phi_{\mu\nu}-2\Phi^{\mu\tau}\nabla_{\tau}\nabla^{\rho}\Phi_{\mu\rho}-\Phi^{\mu}{}_{\mu}\,\Box\,\Phi^{\nu}{}_{\nu}+2\Phi^{\rho}{}_{\rho}\,\nabla_{\mu}\nabla_{\nu}\,\Phi^{\mu\nu}$
$+$ $\frac{1}{L^{2}}(2\,\Phi^{\mu\nu}\Phi_{\mu\nu}+(d-3)\Phi^{\mu}{}_{\mu}\Phi^{\nu}{}_{\nu})+\frac{1}{d-2}(\Phi\,\Box\,\Phi-\frac{1}{L^{2}}(6-2d)\Phi\Phi)$
${\cal L}_{f.}$ $=$ $i\bar{\tilde{\Psi}}^{\mu}\gamma^{\nu}\nabla_{\nu}\tilde{\Psi}_{\mu}-i\bar{\tilde{\Psi}}^{\mu}\nabla_{\mu}\gamma^{\nu}\tilde{\Psi}_{\nu}-i\bar{\tilde{\Psi}}^{\mu}\gamma_{\mu}\nabla_{\nu}\tilde{\Psi}^{\nu}+i(\bar{\tilde{\Psi}}^{\mu}\gamma_{\mu})\gamma^{\nu}\nabla_{\nu}(\gamma^{\rho}\tilde{\Psi}_{\rho})$
$+$ $\frac{i}{2L}((d-2){\bar{\tilde{\Psi}}}^{\mu}{\tilde{\Psi}}_{\mu}-d\bar{\tilde{\Psi}}^{\mu}\gamma_{\mu}\gamma^{\nu}{\tilde{\Psi}}_{\nu})$
$+$ $\frac{i}{d-2}(\bar{\tilde{\Psi}}\gamma^{\mu}\nabla_{\mu}\tilde{\Psi}+\frac{(d-4)}{2L}\bar{\tilde{\Psi}}\tilde{\Psi})$
For the antisymmetric field ${\cal B}_{\mu\nu}$ we get
${\cal L}_{bos.}^{\prime\prime}={\cal B}^{\mu\nu}\,\Box\,{\cal B}_{\mu\nu}-2{\cal B}^{\mu\nu}\,\nabla_{\nu}\nabla^{\tau}\,{\cal B}_{\mu\tau}+\frac{2(d-2)}{L^{2}}{\cal B}^{\mu\nu}{\cal B}_{\mu\nu}=-\frac{1}{3}H^{\mu_{1}\mu_{2}\mu_{3}}H_{\mu_{1}\mu_{2}\mu_{3}}$ (3.15)
where $H_{\mu_{1}\mu_{2}\mu_{3}}=\partial_{[\mu_{1}}{\cal B}_{\mu_{2}\mu_{3}]}$. In case of $d=4$, one has another option for $AdS$ deformation. Namely, one can first dualize the field ${\cal B}_{\mu\nu}$ to a scalar $a(x)$ as $\partial_{\mu}a\sim\epsilon_{\mu\nu\rho\sigma}H^{\nu\rho\sigma}$ and then perform the $AdS_{4}$ deformation to a conformal scalar, with the mass–like term being the same as in (3.2) for the scalar field $\Phi$. As a result, one obtains either the linear ${\cal N}=1$ supermultiplet [40][41], or the Wess–Zumino supermultiplet, where the spin $\frac{1}{2}$ field being $\tilde{\Psi}$. The fields ${\Phi}_{\mu\nu}$ and $\tilde{\Psi}_{\mu}$, in turn form a supermultiplet of the ${\cal N}=1$ supergravity on $AdS_{4}$, constructed in[42].
As an alternative option one can perform the $AdS_{d}$ deformation of the Lagrangian (2) from the beginning
${\cal L}_{bos.}$ $=$ $\phi^{\mu,\nu}\Box\phi_{\mu,\nu}-B\Box A-A\Box B$
-- $2E^{\mu}\nabla_{\mu}B-2C^{\nu}\nabla^{\mu}\phi_{\mu,\nu}-2C^{\nu}\nabla_{\nu}A-2E^{\mu}\nabla^{\nu}\phi_{\mu,\nu}$
-- $C^{\nu}C_{\nu}-E^{\mu}E_{\mu}\,-\frac{1}{L^{2}}\Big{(}-(d-1)\phi^{\mu,\nu}\phi_{\mu,\nu}+(d-3)\phi^{\mu,\nu}\phi_{\nu,\mu}$
-- $2\phi^{\mu}{}_{,\mu}\phi^{\nu}{}_{,\nu}-4\phi^{\mu}{}_{,\mu}(A+B)-(3+d)(A+B)^{2}\Big{)}$
Using the commutation relations given in the Appendix, one can check, that Lagrangian (3.2) is invariant under the gauge transformations
$\delta\phi_{\mu,\nu}=\nabla_{\mu}\lambda_{\nu}+\nabla_{\nu}\rho_{\mu},$
$\delta A=-\nabla^{\nu}\rho_{\nu}-\tau,$
$\delta B=-\nabla^{\nu}\lambda_{\nu}+\tau,$
$\delta C_{\nu}=-\Box\lambda_{\nu}+\frac{1}{L^{2}}(d-1)\rho_{\mu}+\nabla_{\nu}\tau,$ (3.17)
$\delta E_{\mu}=-\Box\rho_{\mu}+\frac{1}{L^{2}}(d-1)\lambda_{\mu}-\nabla_{\mu}\tau.$
Further steps are exactly the same as in the case of a flat background and lead to the Lagrangians (3.2) and (3.15). An analogous procedure, i.e., the diagonalization of the fermionic triplets on $AdS_{d}$ to the spins $\frac{3}{2}$ and $\frac{1}{2}$ has been performed in [39], where further details can be found.
4 Diagonalization for $n=2$
The diagonalization procedure for the case $n=2$ essentially repeats the same steps as for $n=1$. However it has some novel features, such as a presence of the simplest mixed symmetry fields.
One starts from the Lagrangian (2) which in this case reads
${\cal L}_{bos.}$ $=$ $\phi^{\mu_{1}\mu_{2},\nu}\,\Box\,\phi_{\mu_{1}\mu_{2},\nu}-\phi^{\mu_{1}\mu_{2},\nu_{1}}\partial_{\nu_{1}}\partial^{\nu_{2}}\phi_{\mu_{1}\mu_{2},\nu_{2}}-2\phi^{\mu_{1}\mu_{2},\nu}\partial_{\mu_{2}}\partial^{\mu_{3}}\phi_{\mu_{1}\mu_{3},\nu}$
-- $4\phi^{\mu_{1}\mu_{2},\nu}\partial_{\nu}\partial_{\mu_{1}}M_{\mu_{2}}+4\phi^{\mu_{1}\mu_{2},\nu}\partial_{\mu_{1}}\partial_{\mu_{2}}D_{\nu}$
-- $4D^{\nu}\,\Box\,D_{\nu}+2D^{\nu_{1}}\partial_{\nu_{1}}\partial_{\nu_{2}}D^{\nu_{2}}+4D^{\nu}\partial_{\nu}\partial_{\mu}M^{\mu}-2M^{\mu}\Box M_{\mu}$
Next, one decomposes the field $\phi_{\mu_{1}\mu_{2},\nu}$ into a sum
$\phi_{\mu_{1}\mu_{2},\nu}=\tilde{\phi}_{\mu_{1}\mu_{2}\nu}+{\cal B}_{\mu_{1}\mu_{2},\nu}$ (4.2)
where the field $\tilde{\phi}_{\mu_{1}\mu_{2}\nu}(x)$ is totally symmetric
$\tilde{\phi}_{\mu_{1}\mu_{2}\nu}=\frac{1}{3}(\phi_{\mu_{1}\mu_{2},\nu}+\phi_{\mu_{1}\nu,\mu_{2}}+\phi_{\mu_{2}\nu,\mu_{2}})$ (4.3)
and the field ${\cal B}_{\mu_{1}\mu_{2},\nu}$ has the mixed symmetry of the type $Y(2,1)$
${\cal B}_{\mu_{1}\mu_{2},\nu}=\frac{1}{3}(2\phi_{\mu_{1}\mu_{2},\nu}-\phi_{\mu_{1}\nu,\mu_{2}}-\phi_{\mu_{2}\nu,\mu_{2}}),\quad{\cal B}_{\mu_{1}\mu_{2},\nu}={\cal B}_{\mu_{2}\mu_{1},\nu},\quad{\cal B}_{(\mu_{1}\mu_{2},\nu)}=0$ (4.4)
Introducing the linear combinations
$\tilde{D}_{\mu}=\frac{1}{3}(D_{\mu}-M_{\mu}),\quad{\tilde{M}}_{\mu}=M_{\mu}+2D_{\mu}$ (4.5)
and inserting them together with (4.2) into the Lagrangian (4), one finds that it splits into two components (3.3). The first component ${\cal L}_{bos.}^{\prime}$ describes simultaneously spin $3$ and spin $1$ fields and is given by the Lagrangian
${\cal L}_{bos.}^{\prime}$ $=$ $\tilde{\phi}^{\mu_{1}\mu_{2}\mu_{3}}\,\Box\,\tilde{\phi}_{\mu_{1}\mu_{2}\mu_{3}}-3\tilde{\phi}^{\mu_{1}\mu_{2}\nu_{1}}\partial_{\nu_{1}}\partial^{\nu_{2}}\tilde{\phi}^{\mu_{1}\mu_{2}\nu_{2}}$
-- $12\tilde{\phi}^{\mu_{1}\mu_{2}\mu_{3}}\partial_{\mu_{1}}\partial_{\mu_{2}}\tilde{D}_{\mu_{3}}+12\tilde{D}^{\mu}\,\Box\,\tilde{D}_{\mu}+6\tilde{D}^{\mu_{1}}\partial_{\mu_{1}}\partial_{\mu_{2}}\tilde{D}^{\mu_{2}}$
being invariant under the gauge transformations with a symmetric parameter ${\tilde{\Lambda}}_{\mu\nu}$
$\delta\tilde{\phi}_{\mu_{1}\mu_{2}\mu_{3}}=\partial_{(\mu_{1}}{\tilde{\Lambda}}_{\mu_{2}\mu_{3})},\quad\delta\tilde{D}_{\mu}=\partial^{\nu}\tilde{\Lambda}_{\mu\nu},$ (4.7)
The second term in (3.3) describes a reducible representation of the Poincaré group with the mixed symmetry of the type $Y(2,1)$
${\cal L}_{bos.}^{\prime\prime}$ $=$ ${\cal B}^{\mu_{1}\mu_{2},\nu}\,\Box\,{\cal B}_{\mu_{1}\mu_{2},\nu}-{\cal B}^{\mu_{1}\mu_{2},\nu_{1}}\partial_{\nu_{1}}\partial^{\nu_{2}}{\cal B}_{\mu_{1}\mu_{2},\nu_{2}}-2{\cal B}^{\mu_{1}\mu_{2},\nu}\partial_{\mu_{2}}\partial^{\mu_{3}}{\cal B}_{\mu_{1}\mu_{3},\nu}$ (4.8)
-- $4{\cal B}^{\mu_{1}\mu_{2},\nu}\partial_{\mu_{1}}\partial_{\nu}{\tilde{M}}_{\mu_{2}}-\frac{2}{3}{\tilde{M}}^{\mu}\,\Box\,{\tilde{M}}_{\mu}+\frac{2}{3}{\tilde{M}}^{\mu}\,\partial_{\mu}\partial_{\nu}\,{\tilde{M}}^{\nu}$
and is invariant under transformations
$\delta{\cal B}_{\mu_{1}\mu_{2},\nu}$ $=$ $\partial_{(\mu_{1}}\Xi_{\mu_{2})\nu}+\frac{2}{3}\partial_{\nu}{\Lambda}_{\mu_{1}\mu_{2}}-\frac{1}{3}\partial_{(\mu_{1}}\Lambda_{\mu_{2})\nu}$ (4.9)
$\delta{\tilde{M}}_{\mu}$ $=$ $-\partial^{\nu}\Lambda_{\nu\mu}+3\partial^{\nu}\Xi_{\nu\mu}+3\partial_{\mu}\sigma$ (4.10)
with one symmetric $\Lambda_{\mu\nu}$ and one antisymmetric $\Xi_{\mu\nu}$ parameters of gauge transformations. The new parameters are related to the original ones given in (2.7), as
$\tilde{\Lambda}_{\mu\nu}=\frac{1}{3}(\rho_{\mu\nu}+\lambda_{(\mu,\nu)}),\quad\Lambda_{\mu\nu}=\rho_{\mu\nu}-\frac{1}{2}\lambda_{(\mu,\nu)},\quad\Xi_{\mu\nu}=\frac{1}{2}(\lambda_{\mu,\nu}-\lambda_{\nu,\mu})$ (4.11)
In order to decompose the Lagrangian (4) into a sum of the Lagrangians for the individual Fronsdal modes one introduces the fields $\Phi_{\mu_{1}\mu_{2}\mu_{3}}$ and $\Phi_{\mu}$ as
$\tilde{\phi}_{\mu_{1}\mu_{2}\mu_{3}}=\Phi_{\mu_{1}\mu_{2}\mu_{3}}+\frac{1}{d}\eta_{(\mu_{1}\mu_{2}}\Phi_{\mu_{3})},\quad\Phi_{\mu}=\tilde{\phi}^{\nu}{}_{\nu\mu}-2\tilde{D}_{\mu}$ (4.12)
As it was in the case of $n=1$, the precise from of this decomposition is obtained from the requirement that the Lagrangian splits into a sum of Lagrangians for individual Fronsdal modes. However one can get the same result by requiring that the fields $\Phi_{\mu_{1}\mu_{2}\mu_{3}}$ and $\Phi_{\mu}$ have “proper” gauge transformation rules i.e.,
$\delta\Phi_{\mu_{1}\mu_{2}\mu_{3}}=\partial_{(\mu_{1}}{\tilde{\Lambda}}_{\mu_{2},\mu_{3})},\quad\delta{\Phi}_{\mu}=\partial_{\mu}\tilde{\Lambda}$ (4.13)
where the parameter $\tilde{\Lambda}_{\mu\nu}$ is traceless.
For system describing mixed symmetry fields one perform the following field redefinitions
${\cal B}_{\mu_{1}\mu_{2},\nu}=T_{\mu_{1}\mu_{2},\nu}+\frac{1}{2d-6}(2\eta_{\mu_{1}\mu_{2}}T_{\nu}-\eta_{\nu(\mu_{1}}T_{\mu_{2})}),\qquad T_{\nu}={\cal B}^{\mu}{}_{\mu,\nu}-\frac{2}{3}{\tilde{M}}_{\nu}$ (4.14)
This transformation results into a decomposition of the Lagrangian (4.8) into a sum of a standard Lagrangian for a massless vector field $T_{\mu}$
${\cal L}^{s=1}=\frac{3}{2d-6}(T^{\mu}\,\Box\,T_{\mu}-T^{\mu}\partial_{\mu}\partial_{\nu}T^{\nu})$ (4.15)
being invariant under
$\delta T_{\mu}=\partial_{\mu}\xi,\quad\xi=\frac{2}{3}(\Lambda^{\mu}{}_{\mu}-3\sigma)$ (4.16)
and of the Lagrangian for the mixed symmetry field [43]
${\cal L}^{(2,1)}$ $=$ $T^{\mu_{1}\mu_{2},\nu}\,\Box\,T_{\mu_{1}\mu_{2},\nu}-T^{\mu_{1}\mu_{2},\nu_{1}}\partial_{\nu_{1}}\partial^{\nu_{2}}T_{\mu_{1}\mu_{2},\nu_{2}}-2T^{\mu_{1}\mu_{2},\nu}\partial_{\mu_{2}}\partial^{\mu_{3}}T_{\mu_{1}\mu_{3},\nu}$ (4.17)
$+$ $3T^{\mu}{}_{\mu,\nu}\partial_{\mu_{1}}\partial_{\mu_{2}}T^{\mu_{1}\mu_{2},\nu}-\frac{3}{2}T^{\mu}{}_{\mu,\nu}\Box T^{\rho}{}_{\rho,}{}^{\nu}+\frac{3}{2}T^{\mu}{}_{\mu,\nu_{1}}\partial^{\nu_{1}}\partial^{\nu_{2}}T^{\rho}{}_{\rho,\nu_{2}}$
invariant under the same transformations as (4.9), where the symmetric parameter is replaced with
$\Lambda_{\mu\nu}\rightarrow\Lambda_{\mu\nu}+\frac{1}{6-2d}\eta_{\mu\nu}\xi$ (4.18)
In the fermionic sector the Lagrangian (2)
${\cal L}_{f.}$ $=$ $i\,\bar{\Psi}^{\mu_{1}\mu_{2}}\gamma^{\nu}\partial_{\nu}\Psi_{\mu_{1}\mu_{2}}+2i\,\bar{\Psi}^{\mu_{1}\mu_{2}}\partial_{\mu_{1}}\chi_{\mu_{2}}-2i\,\bar{\chi}_{\mu_{1}}\partial_{\mu_{2}}\Psi^{\mu_{1}\mu_{2}}-2i\,\bar{\Sigma}\,\gamma^{\nu}\partial_{\nu}\Sigma$ (4.19)
-- $2i\,\bar{\chi}^{\mu}\gamma^{\nu}\partial_{\nu}\chi_{\mu}+2i\,\bar{\chi}^{\mu}\partial_{\mu}\Sigma-2i\,\bar{\Sigma}\,\partial_{\mu}\chi^{\mu}$
is invariant under gauge transformations
$\delta\Psi_{\mu_{1}\mu_{2}}=\partial_{\mu_{1}}\xi_{\mu_{2}}+\partial_{\mu_{2}}\xi_{\mu_{1}},\quad\delta\chi_{\mu}=-\gamma^{\nu}\partial_{\nu}\xi_{\mu},\quad\delta\Sigma=\partial_{\mu}\xi^{\mu}$ (4.20)
This Lagrangian can be brought into a sum of Lagrangians, each describing an individual Fang-Fronsdal mode by using the transformations
$\tilde{\Psi}_{\mu_{1}\mu_{2}}$ $=$ $\Psi_{\mu_{1}\mu_{2}}-\frac{1}{d}(\gamma_{(\mu_{1}}\gamma^{\mu}\Psi_{\mu_{2})\mu}+\gamma_{(\mu_{1}}\chi_{\mu_{2})})-\frac{1}{d}\eta_{\mu_{1}\mu_{2}}(\Psi^{\mu}{}_{\mu}-2\Sigma)$ (4.21)
$\tilde{\Psi}_{\mu}$ $=$ $\gamma^{\nu}\Psi_{\mu\nu}+\chi_{\mu}-\frac{1}{d-2}\gamma_{\mu}(\Psi^{\nu}{}_{\nu}-2\Sigma),$ (4.22)
$\tilde{\Psi}$ $=$ $\Psi^{\mu}{}_{\mu}-2\Sigma$ (4.23)
Similarly to the case of bosonic fields, the form of these transformations can be obtained from the requirement that the new fields transform “properly” under gauge transformations. Namely, the expression (4.23) can be obtained from the requirement that $\tilde{\Psi}$ is the gauge invariant spin $\frac{1}{2}$ field. For the spin $\frac{5}{2}$ and $\frac{3}{2}$ fields one has the gauge transformation rules
$\delta\tilde{\Psi}_{\mu_{1}\mu_{2}}=\partial_{(\mu_{1}}\tilde{\xi}_{\mu_{2})},\quad\delta\tilde{\Psi}_{\mu}=\partial_{\mu}\tilde{\xi},\quad\tilde{\xi}_{\mu}=\xi_{\mu}-\frac{1}{d}\gamma_{\mu}\gamma^{\nu}\xi_{\nu},\quad\tilde{\xi}=\gamma^{\mu}\xi_{\mu}$ (4.24)
being satisfied by the transformations (4.21) and (4.22). Obviously, these considerations can not determine the numerical coefficients in front of $\Psi$ in the last terms of the equations (4.21) and (4.22), since $\Psi$ is gauge invariant. The only possibility to determine these coefficients is to require that the Lagrangian (4.19) when written in terms of the irreducible modes, splits into a direct sum for individual irreducible representations of the Poincaré group. Noticing that there are always the same combinations of the “untilded” fields in (4.21)–(4.23), it is straightforward to invert these transformations by starting from the bottom one to obtain
$\Psi_{\mu_{1}\mu_{2}}$ $=$ $\tilde{\Psi}_{\mu_{1}\mu_{2}}+\frac{1}{d}\gamma_{(\mu_{1}}\tilde{\Psi}_{\mu_{2})}+\frac{1}{d-2}\eta_{\mu_{1}\mu_{2}}\tilde{\Psi}$ (4.25)
$\chi_{\mu}$ $=$ $-\gamma^{\nu}\tilde{\Psi}_{\nu\mu}+\frac{1}{d}(\gamma_{\mu}\gamma^{\nu}\chi_{\nu}-2\chi_{\mu})$
$\Sigma$ $=$ $\frac{1}{2}\tilde{\Psi}^{\mu}{}_{\mu}+\frac{1}{d}\gamma^{\mu}\tilde{\Psi}_{\mu}+\frac{1}{d-2}\tilde{\Psi}$
Inserting these expressions into the Lagrangian (4.19), one obtains a direct sum of Fang-Fronsdal Lagrangians for spins $\frac{5}{2}$, $\frac{3}{2}$ and $\frac{1}{2}$ with overall factors of $1$, $\frac{1}{d}$ and $\frac{1}{2(d-2)}$.
Finally, using the field redefinitions given in the present Section, and supersymmetry transformations
$\delta\Psi_{\mu_{1}\mu_{2}}=-\gamma^{\rho\nu}\,\epsilon\,\partial_{\rho}\phi_{\mu_{1}\mu_{2},\nu}+\epsilon\,\partial_{(\mu_{1}}B_{\mu_{2})},$ (4.26)
$\delta\Sigma=-\gamma^{\rho\nu}\,\epsilon\,\partial_{\rho}D_{\nu}+\epsilon\partial^{\mu}B_{\mu},$
$\delta\chi_{\mu}=\gamma^{\nu}\,\epsilon\,(\partial^{\mu_{1}}\phi_{\mu\mu_{1},\nu}-\partial_{\mu}D_{\nu})+\gamma^{\nu}\,\epsilon\partial_{\nu}A_{\mu}$
$\delta\phi_{\mu_{1},\mu_{2},\nu}=i\,\bar{\Psi}_{\mu_{1}\mu_{2}}\gamma_{\nu}\,\epsilon,\quad\delta D_{\nu}=i\,\bar{\Sigma}\gamma_{\nu}\,\epsilon,\quad\delta B_{\mu}=-i\,\bar{\chi}_{\mu}\,\epsilon,\quad\delta A_{\mu}=0.$
it is straightforward to obtain supersymmetry transformations at each step described above.
Let us make a few comments on $AdS_{d}$ deformation for the case of $n=2$. Recall, that according to the so-called BMV conjecture [43], performing an $AdS_{d}$ deformation of the Lagrangian (4.17) one loses the gauge invariance with respect to the symmetric parameter $\Lambda_{\mu\nu}$ and as the result the number of on-shell degrees of freedom changes [^5] See [44][46] for a proof of the BMV conjecture, [47] for a discussion for the case of the de Sitter background and [48] for recent progress in the BRST approach.. In order to avoid this issue, one introduces an extra Stueckelberg-type rank-two field, which couples to the original mixed symmetry field in the Lagrangian. This new field transforms under the parameter $\Lambda_{\mu\nu}$ and also brings “ its own” gauge symmetry with a parameter $\xi_{\mu}$. Then one can prove, that after fixing all gauge symmetries, the modified system contains the same degrees of freedom as the original Lagrangian on the flat background. From the consideration above one can conclude, that in order to repeat the procedure described in Subsection 3.2 for the case at hand, one has to follow the option when one performs the splitting of the original Lagrangian into irreducible representations prior to the $AdS_{d}$ deformation. In this case a coupling of the Stueckelberg-like field to the symmetric part of $\phi_{\mu_{1}\mu_{2},\nu}$ can be avoided, and this keeps a possibility for the diagonalization procedure also on $AdS_{d}$. A detailed study of this question, is beyond the scope of the present paper and we leave it for a future work.
5 Diagonalization for a general $n$
The procedure that was shown in detail in the previous two Sections can be generalized to the case of an arbitrary $n$. Using the language of auxiliary oscillators, described in the Section 2, one starts with the Lagrangian (2) and splits the field $|\phi\rangle$ into a sum
$|\phi\rangle=|\tilde{\phi}\rangle+|{\cal B}\rangle$ (5.1)
where the vectors $|\tilde{\phi}\rangle$ and $|{\cal B}\rangle$ contain fields with symmetry of $Y(n+1,0)$ and $Y(n,1)$, respectively
$|\tilde{\phi}\rangle=\frac{1}{(n+1)!}\tilde{\phi}_{\mu_{1},..,\mu_{n}\mu_{n+1}}(x)\,\alpha^{\mu_{1},+}...\alpha^{\mu_{n},+}\psi^{\mu_{n+1},+}|0\rangle$ (5.2)
$|{\cal B}\rangle=\frac{1}{n!}{\cal B}_{\mu_{1},..,\mu_{n},\nu}(x)\,\alpha^{\mu_{1},+}...\alpha^{\mu_{n},+}\psi^{\nu,+}|0\rangle,$
The form of the coupling of the field $|\phi\rangle$ to the fields $|D\rangle$ and $|M\rangle$ in the Lagrangian (2) suggests the field redefinitions
$|\tilde{D}\rangle=-|M\rangle+t|D\rangle,\quad|\tilde{M}\rangle=t^{+}|M\rangle+2|D\rangle$ (5.3)
where the vectors $|\tilde{D}\rangle$ and $|\tilde{M}\rangle$ have the form similar to (5.2), namely
$|\tilde{D}\rangle=\frac{1}{(n-1)!}\tilde{D}_{\mu_{1},..,\mu_{n-1}}(x)\,\alpha^{\mu_{1},+}...\alpha^{\mu_{n-1},+}|0\rangle$ (5.4)
$|{\tilde{M}}\rangle=\frac{1}{(n-2)!}{\tilde{M}}_{\mu_{1},..,\mu_{n-2},\nu}(x)\,\alpha^{\mu_{1},+}...\alpha^{\mu_{n-2},+}\psi^{\nu,+}|0\rangle$
Inverting these transformations
$|D\rangle=\frac{1}{2}(1-\frac{1}{N_{\alpha,\psi}+2}t^{+}t)|\tilde{M}\rangle+t^{+}|\tilde{D}\rangle,\quad|M\rangle=\frac{1}{N_{\alpha,\psi}+2}t|\tilde{M}\rangle-2|\tilde{D}\rangle$ (5.5)
one can check that in terms of the newly introduced fields the Lagrangian has a form (3.3), where
${\cal L}^{\prime}_{bos.}$ $=$ $-\langle\tilde{\phi}|l_{0}-l^{+}l-g^{+}g|\tilde{\phi}\rangle-\langle\phi|l^{+}g^{+}|\tilde{D}\rangle-\langle\tilde{D}|g\,l|\tilde{\phi}\rangle$
$+$ $\langle\tilde{D}|(2l_{0}+l^{+}l)(N_{\alpha,\psi}+2)|\tilde{D}\rangle$
which describes reducible representations of type $Y(n+1,0)$ and the Lagrangian
${\cal L}_{bos.}^{\prime\prime}$ $=$ $-\langle{\cal B}|l_{0}-l^{+}l-g^{+}g|{\cal B}\rangle-\frac{1}{2}\langle{\cal B}|l^{+}\,l^{+}|\tilde{M}\rangle-\frac{1}{2}\langle\tilde{M}|l\,l|{\cal B}\rangle$
$+$ $\frac{1}{4}\langle\tilde{M}|2\,l_{0}(1+\frac{1}{N_{\alpha,\psi}+2}t^{+}t)-{\cal T}(-l^{+}l+g^{+}g){\cal T}$
$+$ $\frac{2}{N_{\alpha,\psi}+2}({\cal T}g^{+}lt+t^{+}l^{+}g{\cal T})|\tilde{M}\rangle,$
where
$N_{\alpha,\psi}=\alpha^{+}\cdot\alpha+\psi^{+}\cdot\psi,\quad{\cal T}=1-\frac{1}{N_{\alpha,\psi}+2}t^{+}t$ (5.8)
which describes reducible representations with mixed symmetry of type $Y(n,1)$. Therefore, the entire system is described as a sum of the bosonic Lagrangians (5), (5) and of the fermionic Lagrangian (2).
In order to simplify the subsequent analysis let us notice, that for the case of totally symmetric fields, one can equivalently reformulate the Lagrangian (5) in terms of the fields that depend only on the bosonic oscillators $\alpha^{+}_{\mu}$ as
${\cal L}^{\prime}_{bos.}=-\langle\tilde{\phi}|l_{0}-l^{+}l|\tilde{\phi}\rangle-\langle\phi|l^{+}l^{+}|\tilde{D}\rangle-\langle\tilde{D}|l\,l|\tilde{\phi}\rangle+\langle\tilde{D}|2l_{0}+l^{+}l|\tilde{D}\rangle$ (5.9)
which is invariant under gauge transformations $\delta|\tilde{\phi}\rangle=l^{+}|\tilde{\lambda}\rangle$ and $\delta|\tilde{D}\rangle=l|\tilde{\lambda}\rangle$. Then the third step for the case of totally symmetric bosonic fields goes as follows [27]. One introduces fields
$|{\cal W}_{n}^{p}\rangle=(M_{11})^{p}|\tilde{\phi}_{n}\rangle-p(M_{11})^{p-1}|\tilde{D}_{n-2}\rangle,\quad\delta|{\cal W}_{n}^{p}\rangle=l^{+}(M_{11})^{p-1}|\tilde{\lambda}\rangle$ (5.10)
and takes an ansatz for irreducible fields in a form of an expansion
$|\Phi_{n}\rangle=\sum_{p=0}^{[\frac{n}{2}]}c_{p}\,(M^{+}_{11})^{p}\,|{\cal W}_{n}^{p}\rangle$ (5.11)
The expansion coefficients $c_{q}$ can be found from the requirement of the double traselesness condition on $|\Phi_{n}\rangle$, which should be satisfied by a Fronsdal mode. Alternatively, most of the coefficients $c_{q}$ (except the ones that multiply gauge-invariant combinations $|{\cal W}_{n}^{0}\rangle$) can also be found from the requirement that the Fronsdal field has the “proper” gauge transformations $\delta|\Phi_{n}\rangle=l^{+}|\Lambda_{n-1}\rangle$, with the traceless gauge parameter. In order to check, that the Lagrangian for reducible higher spin fields splits into a sum of the Lagrangians, each describing an irreducible mode, one has to invert the transformations (5.11) and insert the inverted expression into the Lagrangian (5.9). The fact, that this procedure diagonalizes the Lagrangian was explicitly checked up to the spin four, but as it follows from the discussion above, it is natural to expect, that the diagonalization of Lagrangians works for a general value of $n$.
For the fermionic fields one can proceed in an analogous way. Namely, after defining fields
$|{\cal W}_{n}^{2p}\rangle$ $=$ $(M_{11})^{p}\,|\Psi_{n}\rangle-p(M_{11})^{p-1}\,|\Sigma_{n-2}\rangle$ (5.12)
$|{\cal W}_{n}^{2p+1}\rangle$ $=$ $(M_{11})^{p}\,T\,|\Psi_{n}\rangle+(M_{11})^{p}\,|\chi_{n-1}\rangle-p\,(M_{11})^{p-1}\,T\,|\Sigma_{n-2}\rangle$ (5.13)
one looks for a solution for an irreducible mode $|\tilde{\Psi}_{n}\rangle$ in the form of expansion
$|\tilde{\Psi}_{n}\rangle=\sum_{p=0}^{n}c_{p}\,(T^{+})^{p}\,|{\cal W}_{n}^{p}\rangle$ (5.14)
These expressions generalize (3.11) and (4.21)–(4.23) to an arbitrary value of $n$. The coefficients $c_{p}$ can be determined
$c_{2p-1}(n,d)=(-1)^{p}\frac{(d+2(s-p-2))!!}{(d+2(s-2))!!}=c_{2p}(n,d)$ (5.15)
from the requirement, that the field $|\tilde{\Psi}_{n}\rangle$ is triple gamma-traceless, or alternatively from the requirement, that the irreducible mode transforms “properly” under the gauge transformations with the traceless parameter (see (A.22)–(A.23)).
For mixed symmetry fields, the procedure is essentially the same, with the exception that now one has two types of ${\cal W}$, due to the two possible types symmetries contained in the field $|\tilde{M}\rangle$. Bellow we shall illustrate this procedure for the case when the field $|{\cal B}\rangle$ is of the type $Y(3,1)$, a generalization to an arbitrary $n$ is obvious.
First, using the gauge transformation rules (2.6), one can write the gauge transformations for the fields (5.1) and (5.3) as follows
$\delta|{\cal B}\rangle=(l^{+}-g^{+}t)|\Xi\rangle,$ (5.16)
$\delta|{\tilde{M}}\rangle=(3l+(t^{+}t-1)l)|\Xi\rangle+(3g^{+}+l^{+}t^{+})|\sigma\rangle$ (5.17)
where the parameters of gauge transformations are
$|\Xi\rangle=\frac{1}{2}\Xi_{\mu_{1}\mu_{2},\nu}(x)\alpha^{\mu_{1},+}\alpha^{\mu_{2},+}\psi^{\nu,+}|0\rangle,\quad|\sigma\rangle=\sigma_{\mu}(x)\alpha^{\mu,+}|0\rangle$ (5.18)
The parameter $\Xi_{\mu_{1}\mu_{2},\nu}$ is symmmetrical with respect to its first two indices, but has no symmetry with respect to the third index. Notice also that the gauge transformation rule (5.16) is consistent with the symmetry of the field ${\cal B}_{\mu_{1}\mu_{2}\mu_{3},\nu}$ i.e., with the condition $t|{\cal B}\rangle=0$. Next we define in analogy with (5.10)
$|{\cal W}^{(s)}\rangle=-t^{+}tM_{11}|{\cal B}\rangle+\frac{1}{4}t^{+}t|{\tilde{M}}\rangle$ (5.19)
$|{\cal W}^{(as)}\rangle=(-2+t^{+}t)\,(M_{11}|{\cal B}\rangle-\frac{1}{2}|{\tilde{M}}\rangle),\quad t|{\cal W}^{(as)}\rangle=0$ (5.20)
which contain symmetrical and antisymmetrical tensors of the second rank, respectively. These fields are defined in such a way, that similarly to (5.10) their gauge transformations do not contain divergences of the parameters
$\delta|{\cal W}^{(s)}\rangle=(l^{+}t^{+}M_{12}+g^{+}M_{12})|\Xi\rangle+(l^{+}t^{+}+g^{+})|\sigma\rangle$ (5.21)
$\delta|{\cal W}^{(as)}\rangle=(-l^{+}(2M_{11}+t^{+}M_{12})+g^{+}(2tM_{11}+M_{12}))|\Xi\rangle+(-l^{+}t^{+}+g^{+})|\sigma\rangle$ (5.22)
Finally, we are looking for an irreducible field ${\cal T}_{\mu_{1}\mu_{2}\mu_{3},\nu}$ in the form
$|{\cal T}\rangle=|{\cal B}\rangle+\frac{1}{2(d-2)}(2M_{11}^{+}-M_{12}^{+}t)|{\cal W}^{(s)}\rangle+\frac{1}{d}M_{11}^{+}|{\cal W}^{(as)}\rangle$ (5.23)
where the expansion coefficients can be found after recasting the gauge transformations for the fields $|{\cal T}\rangle$ in the form (5.16), where the parameter of gauge transformations is now traceless [33].
6 Conclusion
There are several directions in which our results may be applied. In particular, the diagonalization procedure seems to be essential tool in the construction of supersymmetric Lagrangians which contain reducible higher spin fields on an $AdS_{d}$ background, a problem that we hope to address in a separate publication. It would also be interesting to apply the diagonalization procedure developed in the present paper to supersymmetric models which contain cubic and higher-order interactions [49][52] either on flat or on $AdS_{d}$ backgrounds similar to how it has been done for the case of current-current interactions [27][28].
Another interesting possibility would be to study applications of the diagonalization procedure to quantum supersymmetric higher spin theories, (see [53][54] for recent progress in supersymmetric quantum higher spin theories).
Finally, let us mention that the diagonalization procedure has not been studied for the case of reducible massive higher spin fields [33], even in the absence of supersymmetry. A possible progress in this direction can prove to be useful for the field-theoretic description of interacting spinning back holes, constructed in terms of massive and massless higher spin triplets [55].
Acknowledgments. We would like to thank Sergei Kuzenko and Yasha Neiman for discussions. The work was supported by the Quantum Gravity Unit of the Okinawa Institute of Science and Technology Graduate University (OIST).
Appendix A Lagrangians and the algebra of operators on AdS space
Although in this paper we were primarily concerned with a flat background, here we give the description of massless bosonic and fermionic higher spin fields for a $d$-dimensional anti-de Sitter space, originally given for $d=4$ in [56][57]. The corresponding equations for the Minkowski space-time can be simply recovered by taking the limit $L\rightarrow\infty$, where $L$ is the radius of the $AdS_{d}$ space.
We use the language of an auxiliary Fock space, given in Section 2. The indexes $\mu,\nu,...$ are Einstein indices and $a,b,...$ are tangent space indices. We use the notation $A\cdot B\equiv A^{\mu}B_{\mu}$. We define symmetrization without a factor in the denominator, for example $A_{(\mu\nu)}=A_{\mu\nu}+A_{\nu\mu}$.
The metric is mostly plus. The $AdS_{d}$ covariant derivatives are realized as
$p_{\mu}=-i(\partial_{\mu}+\omega_{\mu}{}^{ab}\alpha_{a}^{+}\alpha_{b}),\qquad[\alpha_{a},\alpha_{b}^{+}]=\eta_{ab}$ (A.1)
$[p_{\mu},p_{\nu}]=\frac{1}{L^{2}}(\alpha_{\mu}^{+}\alpha_{\nu}-\alpha_{\nu}^{+}\alpha_{\mu})+\frac{1}{2L^{2}}\gamma_{\mu\nu}$ (A.2)
where the last term in (A.2) appears when acting on spinors and $\gamma^{\mu\nu}=\frac{1}{2}(\gamma^{\mu}\gamma^{\nu}-\gamma^{\nu}\gamma^{\mu})$. Further, we introduce operators
$l_{0}=g^{\mu\nu}(p_{\mu}p_{\nu}+i\Gamma^{\lambda}_{\mu\nu}p_{\lambda}),\quad T_{0}=\gamma\cdot p$ (A.3)
$l=\alpha\cdot p,\quad l^{+}=\alpha^{+}\cdot p,\quad T=\alpha\cdot\gamma,\quad T^{+}=\alpha^{+}\cdot\gamma$ (A.4)
and
$N_{\alpha}=\alpha^{+}\cdot\alpha,\quad M_{11}^{+}=\frac{1}{2}\alpha^{+}\cdot\alpha^{+},\quad M_{11}=\frac{1}{2}\alpha\cdot\alpha,$ (A.5)
The action of these operators on a state $|\Phi\rangle$ is translated to the action on a symmetric (spinor) tensor field
$|\Phi\rangle=\frac{1}{s!}\Phi_{\mu_{1},...,\mu_{s}}(x)\alpha^{\mu_{1},+}...\alpha^{\mu_{s},+}|0\rangle$ (A.6)
as follows
$l_{0}|\Phi\rangle\rightarrow-\Box\,\Phi_{\mu_{1},...,\mu_{s}}(x),\quad T_{0}|\Phi\rangle\rightarrow-\gamma^{\mu}\nabla_{\mu}\Phi_{\mu_{1},...,\mu_{s}}(x)$ (A.7)
$l|\Phi\rangle\rightarrow-i\nabla^{\mu_{1}}\Phi_{\mu_{1},...,\mu_{s}}(x),\quad l^{+}|\Phi\rangle\rightarrow-i\nabla_{(\mu}\Phi_{\mu_{1},...,\mu_{s})}(x)$ (A.8)
$T|\Phi\rangle\rightarrow\gamma^{\mu_{1}}\Phi_{\mu_{1},...,\mu_{s}}(x),\quad T^{+}|\Phi\rangle\rightarrow\gamma_{(\mu}\Phi_{\mu_{1},...,\mu_{s})}(x)$ (A.9)
$M^{+}_{11}|\Phi\rangle\rightarrow g_{(\mu\nu}\Phi_{\mu_{1},...,\mu_{s})}(x),\quad M_{11}|\Phi\rangle\rightarrow\frac{1}{2}\Phi^{\mu}{}_{\mu,...,\mu_{s}}(x)$ (A.10)
$N_{\alpha}|\Phi\rangle\rightarrow s\Phi_{\mu_{1},...,\mu_{s}}(x)$ (A.11)
where $\nabla_{\mu}$ and $\Box$ are $AdS_{d}$ covariant derivative and a D‘Alembertian, respectively. Using the commutation relations (A.1)–(A.2) it is straightforward to compute
$[l,l^{+}]$ $=$ $l_{0}-\frac{1}{L^{2}}(4M^{+}_{11}M_{11}-N^{2}_{\alpha}+2N_{\alpha}-dN_{\alpha})-\frac{1}{2L^{2}}\alpha^{\mu,+}\alpha^{\nu}\gamma_{\mu\nu}$
$\equiv$ $\tilde{l}_{0}-\frac{1}{2L^{2}}\alpha^{\mu,+}\alpha^{\nu}\gamma_{\mu\nu}$
$\{T_{0},T_{0}\}=2l_{0}+\frac{2}{L^{2}}(T^{+}T-N_{\alpha})+\frac{1}{2L^{2}}(-d^{2}+d)$ (A.13)
$[T_{0},l]=\frac{1}{L^{2}}(2T^{+}M+\frac{1}{2}(1-2N_{\alpha}-d)T)$ (A.14)
as well as
$[l,M^{+}]=l^{+},\quad[N_{\alpha},l]=l,\quad[N_{\alpha},M_{11}]=2M_{11},\quad[M_{11},M^{+}_{11}]=N_{\alpha}+\frac{d}{2}$ (A.15)
$\{T,T\}=4M_{11},\quad[T,l^{+}]=T_{0},\quad\{T,T_{0}\}=2l,\quad[T,M^{+}_{11}]=T^{+}$ (A.16)
$\{T,T^{+}\}=2N_{\alpha}+d,\quad[N_{\alpha},T]=T$ (A.17)
The Lagrangian for an irreducible massless higher spin field $|\Phi\rangle$ in a $d$ dimensional anti-de Sitter space reads [58]
${\cal L}_{bos.}$ $=$ $-\langle\Phi|\tilde{l}_{0}-l^{+}l-2M^{+}_{11}\tilde{l}_{0}M_{11}+M^{+}_{11}l\,l+l^{+}l^{+}M_{11}-M^{+}_{11}l^{+}lM_{11}$
$+$ $\frac{1}{L^{2}}(6-4N_{\alpha}-2d+M^{+}_{11}(10-4N_{\alpha}-2d)M_{11})|\Phi\rangle.$
Using (A) (without the last term in the r.h.s.) and (A.15) one can check, that the Lagrangian is invariant under gauge transformations
$\delta|\Phi\rangle=l^{+}|\Lambda\rangle$ (A.19)
The field $|\Phi\rangle$ and the parameter of gauge transformations $|\lambda\rangle$ obey off-shell constraints
$(M_{11})^{2}|\Phi\rangle=0,\quad M_{11}|\Lambda\rangle=0$ (A.20)
The Lagrangian for a massless half-integer higher spin field on $AdS_{d}$ is [59][60]
${\cal L}_{f.}$ $=$ $-\langle\Psi|T_{0}-l^{+}T-T^{+}l+T^{+}T_{0}T+l^{+}T^{+}M_{11}+M^{+}_{11}Tl-M^{+}_{11}T_{0}M_{11}$
-- $\frac{i}{2L}(2N_{\alpha}-4+d+T^{+}(2-2N_{\alpha}-d)T-$
-- $M^{+}_{11}(2N_{\alpha}+d)M_{11})|\Psi\rangle$
Using (A) — (A.17) one can check that it is invariant under gauge transformations
$\delta|\Psi\rangle=(l^{+}-\frac{i}{2L}T^{+})|\xi\rangle$ (A.22)
with off-shell constraints on the field $|\Psi\rangle$ and on the parameters of gauge transformations $|\xi\rangle$ being
$T^{3}|\Psi\rangle=0,\quad T|\xi\rangle=0$ (A.23)
Finally, when considering mixed symmetry fields on a flat background, in addition to the previously introduced operators (with $p_{\mu}=-i\partial_{\mu}$ and $L\rightarrow\infty$) we also use
$g=\psi\cdot p,\quad g^{+}=\psi^{+}\cdot p$ (A.24)
$N_{\alpha,\psi}=\alpha^{+}\cdot\alpha+\psi^{+}\cdot\psi,\quad\quad M_{12}^{+}=\alpha^{+}\cdot\psi^{+},\quad M_{12}=\alpha\cdot\psi,$ (A.25)
$t^{+}=\psi^{+}\cdot\alpha,\quad t=\alpha^{+}\cdot\psi,\quad$ (A.26)
The states are now either of the form (A.6) or
$|\tilde{\Phi}\rangle=\frac{1}{s!}\tilde{\Phi}_{\mu_{1},...,\mu_{s},\nu}(x)\alpha^{\mu_{1},+}...\alpha^{\mu_{s},+}\psi^{\nu,+}|0\rangle$ (A.27)
The action of the operators (A.3)–(A.5) on states (A.27) are the same as (A.7)–(A.11). We also have
$g|\tilde{\Phi}\rangle\rightarrow-i\partial^{\nu}\tilde{\Phi}_{\mu_{1},...,\mu_{s},\nu}(x),\qquad\,\,\,\,\,g^{+}|\Phi\rangle\rightarrow-i\partial_{\nu}\Phi_{\mu_{1},...,\mu_{s}}(x)$ (A.28)
$t|\tilde{\Phi}\rangle\rightarrow\tilde{\Phi}_{(\mu_{1},...,\mu_{s},\nu)}(x),\qquad\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,t^{+}|\Phi\rangle\rightarrow\Phi_{\mu_{1},...,\nu}(x)$ (A.29)
$M_{12}|\tilde{\Phi}\rangle\rightarrow\tilde{\Phi}^{\nu}{}_{\mu_{2},...,\mu_{s},\nu}(x),\qquad\,\,\,\,\,\,M_{12}^{+}|\Phi\rangle\rightarrow g_{\nu(\mu_{1}}\Phi_{\mu_{2},...,\mu_{s})}(x)$ (A.30)
$N_{\alpha,\psi}|\tilde{\Phi}\rangle\rightarrow(s+1)\tilde{\Phi}_{\mu_{1},...,\mu_{s},\nu}(x),\qquad N_{\alpha,\psi}|\Phi\rangle\rightarrow s\Phi_{\mu_{1},...,\mu_{s}}(x),$ (A.31)